OpenStudy (anonymous):

write an equation of the line that passes through (2 -7) and is parallel to the line x=5

OpenStudy (anonymous):

Ok, so you need a point and the slope to find the equation of a line.

OpenStudy (anonymous):

You have your point.

OpenStudy (anonymous):

i have spent about 45 minutes reviewing my teachers notes ha had us take, and i jsut don't understand this question!

OpenStudy (anonymous):

And if two lines are parallel they have the same slope.

OpenStudy (anonymous):

i would think so

OpenStudy (anonymous):

what is the slope of line x = 5

OpenStudy (anonymous):

So, using the point slope intercept form ( y-0)=m(x-0) do you think you can figure this out?

OpenStudy (anonymous):

Think about what the slope of x=5 should be.

OpenStudy (anonymous):

well now that i know the formula, i could propably figure it out, he didnt give us a equation! thanks

OpenStudy (anonymous):

not equation, formula to use

OpenStudy (anonymous):

Yeah. most people use y=mx+b. But this is so much easier!

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

No problem.